OpenStudy (anonymous):

In the standard equation of a circle (x-h)^2 = r^2, the radius is represented by R && the center is represented by (-h,-k) TRUE or FALSE ?

jimthompson5910 (jim_thompson5910):

Did you mean (x-h)^2 + (y-k)^2 = r^2 ?

OpenStudy (anonymous):

@jim_thompson5910 YES LOL

jimthompson5910 (jim_thompson5910):

The center of (x-h)^2 + (y-k)^2 = r^2 is (h,k) and not (-h,-k). This is something you'll just have to memorize.

OpenStudy (anonymous):

whys that ?

OpenStudy (anonymous):

so its false correct ?

jimthompson5910 (jim_thompson5910):

Why is it (h,k) or why do you have to memorize it?

OpenStudy (anonymous):

why is it (h,k) ?

jimthompson5910 (jim_thompson5910):

Start with x^2+y^2 = r^2. This is circle centered at (0,0) If you replace 'x' with 'x-h', you'll shift the coordinate system or grid h units to the left...which in turn moves the circle h units to the right. The same idea applies to y as well.